# Discovering overlapping communities in multi-layer directed networks

### v1

@@ 6.2 Real data
(a) Heatmap of the estimated mixed memberships for highly mixed export countries.
(b) Heatmap of the estimated mixed memberships for highly mixed import countries.
Figure 3: Estimated memberships of highly mixed countries.
@@
(a) Row (export) communities
(b) Column (import) communities
Figure 5: Row (export) and column (import) communities discovered by CSPDSoS for FAOMTN. Colors represent communities.
@@ 7 Conclusion
Data and code will be made available on request.
@@ Acknowledgements
Qing’s work was sponsored by the Scientific Research Foundation of Chongqing University of Technology (Grant No: 0102240003) and the Natural Science Foundation of Chongqing, China (Grant No: CSTB2023NSCQ-LZX0048).
@@ A.2 Proof of Theorem 1
$\leq\mathrm{max}_{i\in[n]}(\sum_{j\neq i,j\in[n]}|\sum_{l\in[L]}\sum_{m\in[n]}(A_{l}(i,m)A_{l}(j,m)-\Omega_{l}(i,m)\Omega_{l}(j,m))|+\sum_{l\in[L]}\sum_{m\in[n]}\rho^{2}$
$=\mathrm{max}_{i\in[n]}(\sum_{j\neq i,j\in[n]}|\sum_{l\in[L]}\sum_{m\in[n]}(A_{l}(i,m)A_{l}(j,m)-\Omega_{l}(i,m)\Omega_{l}(j,m))|+\rho^{2}nL.$

### v2

@@ 6.2 Real data
Figure 7: Heatmap of the estimated mixed memberships for highly mixed import/export countries.
Figure 8: Heatmap of the estimated mixed memberships for countries in Table 3.
Figure 9: Heatmap of the estimated mixed memberships for countries in Table 4.
@@
Figure 11: Global map of estimated mixed membership intensity in the six import communities for all the 213 countries. For example, the first sub-figure depicts the membership intensity of each country belonging to the first import community $C_{r,1}$ as estimated by our CSPDSoS algorithm. For $i\in[213]$, recall that the membership intensity for country $i$ in import community 1 is represented by $\hat{\Pi}_{r}(i,1)$ The color of country $i$ on the map corresponds to the magnitude of $\hat{\Pi}_{r}(i,1)$, with darker colors indicating higher membership intensity and lighter colors indicating lower membership intensity. Similar arguments for the other five import communities.
Figure 12: Global map of estimated mixed membership intensity in the six export communities for all the 213 countries.
@@ 7 Conclusion
Data will be made available on request. The MATLAB codes for the proposed algorithm are provided in B.
@@ Acknowledgements
Qing’s work was sponsored by the Scientific Research Foundation of Chongqing University of Technology (Grant No. 2024ZDR003), the Science and Technology Research Program of Chongqing Municipal Education Commission (Grant No. KJQN202401168), and the Natural Science Foundation of Chongqing, China (Grant No: CSTB2023NSCQ-LZX0048).
@@ A.2 Proof of Theorem 1
$\leq\mathrm{max}_{i\in[n]}(\sum_{j\neq i,j\in[n]}|\sum_{l\in[L]}\sum_{m\in[n]}(A_{l}(i,m)A_{l}(j,m)-\Omega_{l}(i,m)\Omega_{l}(j,m))|+\sum_{l\in[L]}\sum_{m\in[n]}\rho^{2})$
$=\mathrm{max}_{i\in[n]}(\sum_{j\neq i,j\in[n]}|\sum_{l\in[L]}\sum_{m\in[n]}(A_{l}(i,m)A_{l}(j,m)-\Omega_{l}(i,m)\Omega_{l}(j,m))|)+\rho^{2}nL.$
@@
Appendix B MATLAB codes of CSPDSoS
The MATLAB codes of CSPDSoS are provided below:
function [Pi_r, Pi_c] = CSPDSoS(A_all, K)
% Inputs:
% A_all - A tensor with A_all(:,:,l) being the adjacency matrix of layer l
% K - Number of communities to detect
% Outputs:
% Pi_r - Estimated row membership matrix
% Pi_c - Estimated column membership matrix
% Number of nodes
n = size(A_all, 1);
% Number of layers
L = size(A_all, 3);
% Initialize the debiased aggregation matrices
S_r = zeros(n, n);
S_c = zeros(n, n);
% Compute the debiased aggregation matrices
for l = 1:L
Al = A_all(:, :, l);
Drl = diag(sum(Al, 2)); % Diagonal matrix of row sums
Dcl = diag(sum(Al, 1)); % Diagonal matrix of column sums
S_r = S_r + Al * Al - Drl; % Debiased aggregation matrix for sending patterns
S_c = S_c + Al * Al - Dcl; % Debiased aggregation matrix for receiving patterns
end
% Compute the leading K eigenvectors of S_r and S_c
[U_r, ~] = eigs(S_r, K); % Leading K eigenvectors for sending patterns
[U_c, ~] = eigs(S_c, K); % Leading K eigenvectors for receiving patterns
% Find the indices of the K purest nodes using SPA
[I_r] = SPA(U_r, K); % Indices of pure nodes for sending patterns
[I_c] = SPA(U_c, K); % Indices of pure nodes for receiving patterns
% Compute the membership matrix for sending patterns
C_r = U_r(I_r, :); % Corner matrix in U_r
Y_r = max(0, U_r / C_r);
Pi_r = zeros(n, K); % Initialize row-wise membership matrix
for i = 1:n
Pi_r(i, :) = Y_r(i, :) / sum(Y_r(i, :));
end
% Compute the membership matrix for receiving patterns
C_c = U_c(I_c, :); % Corner matrix in U_c
Y_c = max(0, U_c / C_c);
Pi_c = zeros(n, K); % Initialize column-wise membership matrix
for i = 1:n
Pi_c(i, :) = Y_c(i, :) / sum(Y_c(i, :));
end
% Nested function: SPA (Successive Projection Algorithm)
function [pure] = SPA(X, K)
% SPA: Successive Projection Algorithm
% Inputs:
% X - Feature matrix
% K - Number of purest nodes to select
% Outputs:
% pure - Indices of the K purest nodes
% Initialize empty array for pure node indices
pure = [];
% Compute row norms of the feature matrix
row_norm = vecnorm(X’).^2’;
% Iterate K times to select the K purest nodes
for i = 1:K
% Find the node with the maximum row norm
[~, idx_tmp] = max(row_norm);
% Add the index to the list of pure nodes
pure = [pure idx_tmp];
% Store the eigenvector of the selected pure node
U(i, :) = X(idx_tmp, :);
for j = 1:i-1
U(i, :) = U(i, :) - U(i, :) * (U(j, :)’ * U(j, :));
end
% Normalize the orthogonalized eigenvector
U(i, :) = U(i, :) / norm(U(i, :));
row_norm = row_norm - (X * U(i, :)’).^2;
end
end
end